OpenStudy (anonymous):

How do I solve? How many grams of a stock solution that is 85.5 percent H2SO4 by mass would be needed to make 150 grams of a 45.0 percent by mass solution? Show all of the work needed to solve this problem. Do I use C1V1= C2V2

OpenStudy (abmon98):

C1V1=C2V2 is used when dilution is involved so is in this case we will not use it. you need x gram out of 150 to give a 45% mass solution so 150*45/100=67.5 grams 85.5*x/100=67.5 x represent the amount in grams of stock solution

OpenStudy (somy):

there is another formula connecting mass and concentration C1m1= C2m2 in such type of questions 'percentage' is thought of as concentration so use it same as i told you for C!V1=C2V2

OpenStudy (anonymous):

And M stands for mass right?

OpenStudy (anonymous):

@Somy

OpenStudy (somy):

yes

OpenStudy (anonymous):

Okay so 85.5*? = 45.0*150

OpenStudy (somy):

yes

OpenStudy (anonymous):

6750/85.5?

OpenStudy (somy):

yup)